### v1

# Octax: Accelerated CHIP-8 Arcade Environments for Reinforcement Learning in JAX

Abstract
Reinforcement learning (RL) research requires diverse, challenging environments that are both tractable and scalable. While modern video games may offer rich dynamics, they are computationally expensive and poorly suited for large-scale experimentation due to their CPU-bound execution. We introduce Octax, a high-performance suite of classic arcade game environments implemented in JAX, based on CHIP-8 emulation, a predecessor to Atari, which is widely adopted as a benchmark in RL research. Octax provides the JAX community with a long-awaited end-to-end GPU alternative to the Atari benchmark, offering image-based environments, spanning puzzle, action, and strategy genres, all executable at massive scale on modern GPUs. Our JAX-based implementation achieves orders-of-magnitude speedups over traditional CPU emulators while maintaining perfect fidelity to the original game mechanics. We demonstrate Octax’s capabilities by training RL agents across multiple games, showing significant improvements in training speed and scalability compared to existing solutions. The environment’s modular design enables researchers to easily extend the suite with new games or generate novel environments using large language models, making it an ideal platform for large-scale RL experimentation.
1 Introduction
Modern reinforcement learning (RL) research (Sutton & Barto, 2018) demands extensive experimentation to achieve statistical validity, yet computational constraints severely limit experimental scale. RL papers routinely report results with fewer than five random seeds due to prohibitive training costs (Henderson et al., 2018; Colas et al., 2018; Agarwal et al., 2021; Mathieu et al., 2023; Gardner et al., 2025). While understandable from a practical standpoint, this undersampling undermines statistical reliability and impedes algorithmic progress. Environment execution creates this bottleneck: while deep learning has embraced end-to-end GPU acceleration, RL environments remain predominantly CPU-bound. Originally designed under severe hardware constraints, classic arcade games represent a solution for scalable RL experimentation. The Atari Learning Environment (ALE) (Bellemare et al., 2013) has established itself as a standard RL benchmark, although existing implementations remain fundamentally CPU-bound. As noted by Obando-Ceron & Castro (2020), the Rainbow paper (Hessel et al., 2018) required 34,200 GPU hours (equivalent to 1,425 days) of experiments, a computational cost that is prohibitively high for small research laboratories. In this paper, we propose an alternative approach for training RL agents in environments with mechanisms similar to ALE, with significantly reduced computational cost.
Contributions.
We introduce Octax  [^1] The repository containing all source code, experiments, and data is available at: https://github.com/riiswa/octax, a suite of arcade game environments implemented in JAX (Bradbury et al., 2018a) through CHIP-8 emulation. CHIP-8, a 1970s virtual machine specification contemporary with early Atari systems, became the foundation for numerous classic games spanning puzzle, action, and strategy genres. CHIP-8’s constraint-driven design creates games with similar cognitive demands to Atari while enabling efficient vectorized emulation that scales to thousandsof parallel instances. The JAX ecosystem has rapidly emerged as a solution for scalability in RLresearch but lacks native environments, particularly image-based ones. Our framework addresses this gap by transforming classic games into fully vectorized, GPU-accelerated simulations. Thesesimulations run thousands of game instances in parallel while maintaining perfect fidelity to the original mechanics. This approach dramatically reduces experiment times. Experiments that previously required days or weeks can now be completed in hours. This efficiency makes comprehensive hyperparameter sweeps and ablation studies computationally feasible. The modular design facilitates extension with new games or automated generation using large language models that can directly output CHIP-8 assembly code. Figure 1 provides an overview of the integrated CHIP-8 games.
Outline.
First, we present the our end-to-end JAX implementation of classic arcade environments through CHIP-8 emulation (Section 3). Second, we demonstrate diverse learning dynamics through PPO evaluation across 16 games (Section 4.1). Third, we achieve 350,000 environment steps per second (1.4 million frames per second) on consumer-grade hardware, substantially outperforming CPU-based solutions (Section 4.2). Fourth, we establish an LLM-assisted pipeline for automated environment generation that creates meaningful difficulty gradients (Section 4.3).
Figure 1: Overview of CHIP-8 game environments implemented in Octax.
2 Related Work
Game environments have proven essential for RL research because they provide engaging, human-relevant challenges with clear success metrics. The Arcade Learning Environment (ALE) Bellemare et al. (2013) demonstrated this principle by establishing Atari 2600 games as the standard RL benchmark, enabling breakthrough algorithms like DQN (Mnih et al., 2015) and Rainbow (Hessel et al., 2018). The success of these classic arcade games stems from their constraint-driven design: simple rules that yield complex behaviors, deterministic dynamics that enable reproducible experiments, and visual complexity that tests spatial reasoning without overwhelming computational resources.
While algorithmic advances demand increasingly large-scale experiments with thousands of parallel environments and extensive hyperparameter sweeps, traditional game environments remain CPU-bound and poorly suited for parallel execution. This mismatch has driven a progression of solutions, each addressing different aspects of the scalability problem.
Game-based RL environment platforms. Increasingly sophisticated gaming platforms have been developed to test different dimensions of learning performance. NetHack Learning Environment (Küttler et al., 2020) provides procedurally generated roguelike challenges that test long-term planning, while Crafter (Hafner, 2021) offers simplified Minecraft-like environments focused on resource management. These environments expand cognitive challenges beyond arcade games, but their CPU-based implementations compound the scalability problem.
CPU high-performance solutions. Several projects have focused on optimizing CPU-based environment execution. EnvPool (Weng et al., 2022) achieves substantial speed improvements through highly optimized C++ implementation, demonstrating up to 1 million Atari frames per second on high-end hardware. PufferLib (Suarez, 2025) provides environments written entirely in C, achieving millions of steps per second through over 20,000 lines of optimized code. While these approaches improve CPU throughput, they retain fundamental limitations: costly CPU-GPU data transfers during training and require C implementation in a Python-dominated field.
GPU-accelerated RL environments. GPU-accelerated solutions target the constraint more directly by moving environment execution to accelerators. CUDA Learning Environment (CuLE) (Dalton et al., 2020) provides a pioneering CUDA port of ALE, achieving 40-190 million frames per hour on single GPUs. Isaac Gym (Makoviychuk et al., 2021) demonstrates similar principles for robotics tasks, achieving 2-3 orders of magnitude speedups over CPU approaches by running thousands of environments simultaneously. These GPU approaches solve computational bottlenecks but introduce NVIDIA hardware dependence and substantial per-environment engineering costs.
JAX-based environments. The adoption of JAX (Bradbury et al., 2018b) has enabled natively accelerated environments that combine portability across hardware with end-to-end GPU acceleration. Brax (Freeman et al., 2021) established viability through MuJoCo-like physics simulation, while Gymnax (Lange, 2022) provides JAX implementations of classic control tasks and simplified environments from BSuite (Osband et al., 2019) and MinAtar (Young & Tian, 2019). Specialized environments target specific research needs: XLand-MiniGrid (Nikulin et al., 2024) and Navix (Pignatelli et al., 2024) focus on gridworld navigation, Jumanji (Bonnet et al., 2023) spans domains from simple games to NP-hard combinatorial problems, Pgx (Koyamada et al., 2023) provides classic board games, and PuzzleJAX Earle et al. (2025) enables dynamic compilation of puzzle games.
Despite this coverage, a critical gap remains: classic arcade games. While MinAtar provides simplified versions of Atari games, the full visual complexity and authentic game mechanics of classic arcade games remain absent from the JAX ecosystem. Octax addresses this gap by providing the first end-to-end JAX implementation of classic arcade games through CHIP-8 emulation, delivering computational benefits while preserving the engaging gameplay mechanics that made arcade games valuable for algorithmic development.
3 Octax: The accelerated CHIP-8 Platform
Figure 2: Octax architecture: ROM loading, CHIP-8 emulation pipeline, and RL environment integration. The system transforms game ROMs through fetch-decode-execute cycles into vectorized JAX operations suitable for GPU acceleration.
This section presents our JAX implementation of CHIP-8 emulation. We detail the design decisions that enable GPU acceleration while maintaining behavioral fidelity to original games, and explain how CHIP-8’s architecture provides an optimal foundation for scalable experimentation in RL. Figure 2 summarizes this section.
3.1 Why CHIP-8 for RL research?
CHIP-8 represents a strategic choice for RL environment design. Created in the 1970s as a virtual machine specification, CHIP-8 features a 64×32 monochrome display, 16 registers, 4KB memory, and 35-instruction set. These constraints, originally imposed by early microcomputer limitations, create several research advantages.
The platform provides image-based environments comparable to Atari games while offering some computational advantages. The 4KB memory footprint allows thousands of simultaneous game instances without memory constraints. The simple instruction set reduces emulation overhead compared to complex modern processors. The deterministic execution model ensures experimental reproducibility across different hardware configurations.
The platform supports everything from precise action games requiring split-second timing to complex puzzles demanding long-horizon planning. The 16-key input system provides sufficient complexity for interesting control challenges while remaining tractable for systematic analysis. Most importantly, CHIP-8 games are inherently modifiable and analyzable: their simple assembly code can be automatically generated, modified, and assessed for difficulty, enabling novel research directions in environment design and curriculum learning. This combination of Atari-like visual complexity with modern computational efficiency makes CHIP-8 well-suited for the JAX ecosystem, where extensive parallelization can transform week-long experiments into hour-long runs.
3.2 How does Octax work?
Octax converts CHIP-8 ROMs [^2] ROM stands for Read-Only Memory, a type of storage originally used in game cartridges to hold software that cannot be modified by the user. into vectorized RL environments while maintaining compatibility with original games. The implementation leverages JAX’s functional programming model and vectorization capabilities to enable GPU acceleration.
ROM loading and initialization. Game data is loaded from .ch8 files into the emulator’s 4KB memory space starting at address 0x200, following the standard CHIP-8 program layout first introduced in Weisbecker (1978). The system initializes with font data at address 0x50, sixteen general-purpose registers (V0-VF), an index register (I), a program counter (PC), and the 64×32 monochrome display buffer.
Fetch-decode-execute cycle. The core emulation loop implements the classic processor cycle using JAX primitives. The fetch() function retrieves 16-bit instructions from memory and advances the program counter. The decode() function extracts instruction components through bitwise operations, identifying opcodes, register indices, and immediate values. The execute() function uses JAX’s lax.switch for GPU-compatible instruction dispatch to specialized handlers.
Vectorized instruction execution. Instruction handlers follow JAX’s functional programming model, treating state as immutable and returning updated copies. ALU operations handle arithmetic and bitwise logic with carry/borrow flag management. Control flow instructions implement jumps, calls, and conditional operations using lax.cond. The display system uses vectorized operations to render sprites across the entire framebuffer simultaneously.
Environment integration. The OctaxEnv wrapper transforms the emulator into a standard RL interface. Each RL step executes multiple CHIP-8 instructions to maintain authentic game timing relative to the original 700Hz instruction frequency. The default frame skip setting preserves realistic game dynamics. Observations consist of the 64×32 display with 4-frame stacking, producing (4, 64, 32) boolean arrays. Actions map from discrete RL outputs to game-specific key subsets plus a no-op option. The wrapper manages delay and sound timers at 60Hz and executes startup sequences to bypass menu screens.
3.3 How to transform games into RL environments?
Converting CHIP-8 games into RL environments requires extracting reward signals and termination conditions from game-specific memory layouts and register usage patterns.
Score function design. Games store scores in different registers using various encoding schemes. Octax provides game-specific score_fn functions that extract scores from appropriate memory locations. Brix stores its score in register V5, incrementing with each destroyed brick. Pong encodes scores in BCD format within register V14, requiring score = (V[14] // 10) - (V[14] % 10) to compute player advantage. This flexibility allows researchers to experiment with alternative reward formulations based on different game state components.
Termination logic. Games signal completion through different register states that must be identified through analysis. Brix terminates when lives (V14) reach zero, while Tetris uses a dedicated game-state register (V1) that equals 2 on game over. Some games require compound conditions: Space Flight ends when either lives reach zero or a level completion counter exceeds a threshold, implemented as terminated = (V[9] == 0) | (V[12] >= 0x3E).
Action space optimization. Most games use subsets of the 16-key hexadecimal keypad. Octax supports custom action_set arrays that map RL action indices to relevant keys. Pong requires only keys 1 and 4 for paddle movement, while Worm uses directional keys 2, 4, 6, 8. This reduces action space size and accelerates learning by eliminating irrelevant inputs.
Initialization handling. Many games include menu screens that interfere with RL training. Octax supports startup_instructions parameters that automatically execute instruction sequences during environment reset, bypassing menus to begin gameplay immediately.
We address CHIP-8’s non-standardized scoring and termination by combining static ROM analysis and dynamic memory monitoring during gameplay, as detailed in Appendix C.
3.4 Which games does Octax support?
Octax provides a curated collection of classic CHIP-8 games across multiple genres and difficulty levels. The current implementation includes 21 titles, with additional games planned for future releases. All environments maintain full compatibility with both Gymnasium and Gymnax APIs.
Category Available Games Required Capabilities
Puzzle Tetris, Blinky, Worm Long-horizon planning, spatial reasoning
Action Brix, Pong, Squash, Vertical Brix, Wipe Off, Filter Timing, prediction, reactive control
Strategy Missile Command, Rocket, Submarine, Tank Battle, UFO Resource management, tactical decisions
Exploration Cavern (7 levels), Flight Runner, Space Flight (10 levels), Spacejam! Spatial exploration, continuous navigation
Shooter Airplane, Deep8, Shooting Stars Simple reaction, basic timing
Table 1: Currently implemented games in Octax.
The games (Figure 1) vary across multiple dimensions of difficulty and cognitive demand. Temporal complexity ranges from immediate reactions to long-term planning requirements. Spatial complexity spans single-screen environments to multi-screen worlds requiring navigation. Reward structures include both dense scoring mechanisms and sparse achievement-based systems. This systematic variation enables controlled studies of algorithmic performance across different challenge types while maintaining a unified technical framework for fair comparison. A categorization of these games is provided in Table 1, with more detailed descriptions available in Appendix C.2.
4 Experimental Evaluation
We evaluate Octax through RL training experiments across 16 diverse CHIP-8 games. Our goal is to demonstrate that the environments exhibit varied difficulties and learning dynamics suitable for RL research and benchmark the platform’s computational performance.
4.1 How do RL agents learn in Octax?
Figure 3: PPO learning curves across 16 games: Interquartile Mean (IQM) returns using 10th-90th percentile ranges over 5M timesteps, with confidence intervals computed across 12 random seeds.
We train Proximal Policy Optimization (PPO) (Schulman et al., 2017) agents across our game suite due to its widespread adoption and proven scalability with parallel environments (Rudin et al., 2022).
Network architecture.
Our PPO agent [^3] Based on Rejax implementation (Liesen et al., 2024). uses a convolutional neural network designed for Octax’s (4, 64, 32) stacked observations. The feature extractor consists of three convolutional layers with 32, 64, and 64 filters respectively. These layers use kernel sizes of (8,4), 4, and 3 with corresponding strides of (4,2), 2, and 1. Extracted features are flattened and fed to separate actor and critic heads, each containing a single 256-unit hidden layer with ReLU activation throughout.
Training configuration.
We combine grid search optimization (detailed in Appendix 4) on Pong with CleanRL’s standard Atari PPO hyperparameters (Huang et al., 2022). This yields GAE lambda of 0.95, clipping epsilon of 0.2, value function coefficient of 0.5, and entropy coefficient of 0.01. Each experiment uses 512 parallel environments with 32-step rollouts, 4 training epochs per update, and 32 minibatches for gradient computation. We apply the Adam optimizer (Kingma & Ba, 2014) with learning rate $5\times 10^{-4}$ and gradient clipping for stable training across 5 million timesteps per environment.
Experimental setup.
We conduct 12 independent training runs per game using different random seeds. All experiments run on a single NVIDIA A100 GPU with 24 concurrent training sessions. Agent performance is assessed every 131,072 timesteps on 128 parallel environments.
Results analysis.
The training curves in Figure 3 reveal distinct learning profiles across games. We observe three main patterns that reflect different cognitive demands. Rapid plateau games (Airplane, Brix, Deep, Filter, Blinky) show quick initial learning followed by stable performance, suggesting clear reward signals. Gradual improvement games (Submarine, Tank, UFO) exhibit sustained learning throughout training, indicating either sparser reward structures or more complex strategic requirements. Limited performance games, like Tetris, exhibit significant variance with little absolute progress, making them difficult for standard policy gradient methods. Similarly, in Worm (a Snake clone), agents often manage to eat only a single apple before dying.
These learning profiles support the cognitive diversity of CHIP-8 environments, demonstrating that different games test varied aspects of learning and control. Individual training runs averaged 65 minutes each, with 24 experiments running concurrently, achieving approximately 30,800 environment steps per second across all parallel sessions.
4.2 How does Octax scale with parallelization?
Figure 4: Performance scaling of Octax and EnvPool across parallelization levels. The solid purple line is the number of steps per second (higher is better), and the dashed green line is the total execution time in seconds (lower is better).
Experimental setup.
We measure environment throughput across different parallelization levels to quantify Octax’s computational advantages. This experiment isolates pure computational benefits by fixing the game (Pong) and agent behavior (constant action) while varying parallel environment instances. Since all environments execute identical CHIP-8 computational cycles, these performance measurements apply uniformly across the entire game suite. To better interpret our results, we compare against EnvPool because it is widely adopted in RL research, using ALE Pong to assess CPU vs. GPU-based environment scalability.
Configuration.
We benchmark on a consumer-grade workstation with an RTX 3090 (24GB VRAM), 32GB RAM, and an Intel i7 processor (20 cores). We measure execution time for 100-step rollouts across varying parallel environment counts, with 50 independent measurements per configuration. The primary metric is environment steps per second, calculated as (number of environments × 100 steps) divided by execution time, where each step represents 4 frames due to Octax’s default frame skip setting.
Performance results.
Figure 4 demonstrates near-linear scaling up to 350,000 steps (or 1,4M frames) per second with 8,192 parallel environments before hitting VRAM limitations. EnvPool running ALE Pong with all available CPU cores shows reduced scaling, plateauing around 25,000 steps per second due to CPU saturation. Octax achieves a 14× improvement in computational efficiency at high parallelization levels, reducing the computational cost of large-scale RL experiments. We also measured GPU memory usage across different environment counts, finding that execution memory scales linearly with the number of parallel environments with our benchmark script, consuming approximately 2 MB of GPU memory per environment
4.3 How do LLMs assist environment creation?
Figure 5: Rendering of the Target Shooter game, generated by an LLM, showing the player (left, circular object) and target (right, cross-shaped object).
Figure 6: PPO training performance on LLM-generated environments with varying difficulty.
Large language models (LLMs) have demonstrated a strong capability in code generation across diverse programming languages, enabling the automated creation of environments in RL research. Here we explore Octax’s capacity to accelerate research by leveraging LLMs to generate novel tasks, extending beyond manually designed game suites toward automated environment synthesis, as explored in Faldor et al. (2024).
Context.
During Octax’s development, we encountered several games where reward and termination logic proved difficult to extract through manual analysis of game mechanics. In these cases, we decompiled ROMs to obtain CHIP-8 assembly code and successfully employed LLMs to generate appropriate score_fn and terminated_fn functions by analyzing the assembly instructions. This process revealed LLMs’ capability to understand low-level game logic and translate it into RL-compatible reward structures. This success motivated us to investigate the reverse pipeline: using LLMs to generate complete CHIP-8 games from high-level descriptions, then leveraging Octax’s scalable simulation to evaluate these procedurally created environments.
Figure 7: Environment generation pipeline.
Automated environment generation pipeline. Our pipeline consists of seven replicable steps for automatic CHIP-8 game generation. In Step 1, we construct a corpus of CHIP-8 tutorials, documentation, and programming examples, ensuring the LLM understands the architecture’s instruction set, memory layout, and common coding patterns. In Step 2, we embed this corpus into a prompt (detailed in Appendix E.1) that guides the LLM to produce syntactically correct CHIP-8 programs from high-level instructions. In Step 3, we provide a description of the game with desired mechanics, objectives, and constraints. In Step 4, the LLM generates the initial CHIP-8 code based on the provided description. In Step 5, an automated feedback loop between the LLM and a CHIP-8 compiler iteratively refines the code based on compilation errors until successful. In Step 6, Python wrapper functions for score_fn and terminated_fn are automatically generated, translating CHIP-8 registers into RL-compatible reward and termination signals. Finally, in Step 7, the game description is augmented to increase difficulty or introduce new challenges. Both the new description and the previously generated game are added to the LLM’s context before next iteration. Figure 7 summarizes the automated environment generation pipeline.
Target Shooter case study. We validated this pipeline using Claude Opus 4.1, known for its proficiency in programming, with the following description: "Target Shooter – Targets appear randomly on the screen, and the player moves a crosshair to shoot them. Score increases per hit, and the game ends after a fixed number of targets." The system successfully generated three progressive difficulty levels: static targets for basic aiming skills, time-limited targets introducing decision pressure, and moving targets with time constraints requiring predictive aiming. Each level maintains consistent register mappings for score and termination, simplifying Octax compatibility. Figure 6 shows how the LLM-generated environment visual appearance. All the code generated by the LLM is given in Appendix E.2,
RL experiments. Using identical PPO configurations from Section 4.1, we trained agents on the three generated difficulty levels over 5M timesteps. Figure 6 demonstrates clear performance stratification across difficulty levels: Level 1 agents achieved optimal returns of 10.0 with rapid convergence by 1M timesteps, Level 2 agents plateaued at 9.0 returns with moderate learning speed, while Level 3 agents reached 8.0 returns with the slowest progression. The inverse relationship between difficulty level and both final performance and sample efficiency indicates that our LLM-generated environments successfully create a meaningful difficulty gradient. This proof-of-concept demonstrates the feasibility of automated environment generation for RL research via Octax, with promising applications in curriculum learning, open-endedness, and continual learning scenarios.
5 Conclusion
We introduced Octax, a JAX-based CHIP-8 emulation platform that provides GPU-accelerated arcade game environments for reinforcement learning research. Our implementation achieves significant performance improvements over CPU-based alternatives, enabling experiments with thousands of parallel environments while maintaining perfect behavioral fidelity to original games. Through PPO evaluation across 16 diverse games, we demonstrated varied learning dynamics that highlight the cognitive diversity within classic arcade environments. The platform’s modular design enables both manual game integration and automated environment generation using large language models, providing researchers with flexible experimental design options.
Societal and environmental impact. Octax enables more rigorous evaluation with larger sample sizes, addressing reproducibility concerns that affect institutions with limited computational resources. This implementation can reduce energy consumption compared to resource-intensive benchmarks such as ALE: experiments that once required top-tier clusters can now run efficiently on a single GPU, potentially saving significant compute time and resources.
Limitations. The GPU-based architecture faces performance constraints due to CHIP-8’s variable instruction execution complexity. JAX synchronization across parallel environments means each step’s execution time depends on the slowest instruction among CHIP-8’s 35 operations, typically display rendering or complex ALU operations. The absence of established maximum scores across our game suite prevents the assessment of whether agents approach theoretical performance limits, limiting evaluation of algorithmic performance ceilings.
Future work. Octax can expand through community contributions, with hundreds of compatible ROMs available online. The LLM-assisted environment generation pipeline enables curriculum learning and open-ended research through procedurally generated games that provide task diversity. We plan to investigate emulator optimizations including instruction-level parallelization strategies and adaptive batching to address synchronization bottlenecks from variable execution times. We also aim to extend platform support to Super-CHIP8 and XO-CHIP variants: Super-CHIP8 offers higher resolution displays (128×64) and extended instruction sets originally developed for HP48 calculators, while XO-CHIP provides color graphics, improved audio, and expanded memory while maintaining backward compatibility. These extensions would enable Octax to support more sophisticated games and visual complexity while preserving the computational efficiency advantages of the JAX-native architecture. Many CHIP-8 games feature multi-agent or multi-player mechanics, which we plan to support in future platform releases. The platform’s high-throughput capabilities also position it well for offline RL research, enabling the efficient creation of large-scale datasets and the comprehensive evaluation of offline algorithms across diverse game environments.
Reproducibility Statement
We provide complete resources to ensure reproducibility of our results. The Octax source code, including all 21 game environment implementations, JAX-based CHIP-8 emulator, and training scripts, is available as supplementary material. Our experimental setup uses standard PPO hyperparameters detailed in Section 4.1, with hardware specifications and performance benchmarking configurations provided in Section 4.2. All training experiments use identical network architectures and hyperparameters across games, enabling direct replication of our learning curves in Figure 3. For the LLM-assisted environment generation pipeline in Section 4.3, we include the prompt templates and generated CHIP-8 assembly code in Appendix E. The modular design of Octax allows researchers to extend our game suite using the technical specifications in Section 3. The repository containing all source code, experiments, and data is available at: https://github.com/riiswa/octax.
Appendix A Use of Large Language Models
We used Large Language Models in three capacities during this research. First, Claude Opus 4.1 serves as a core research component in Section 4.3, where we demonstrate automated CHIP-8 game generation from high-level descriptions. This represents a novel research contribution, with all generated code validated through compilation and RL experiments. Second, we employed Claude Sonnet 4 for writing assistance, including text refinement, rephrasing technical concepts, and improving academic tone. Third, LLMs generated code documentation, docstrings, and tutorial content. All research ideas, experimental design, and scientific claims originate from the authors. We did not use LLMs for ideation, hypothesis formation, or result interpretation. We manually reviewed and validated all LLM-assisted content for accuracy and take full responsibility for all presented content.
Appendix B CHIP-8 Technical Specifications
B.1 Platform Overview
CHIP-8 was created by Joseph Weisbecker at RCA in the mid-1970s as a virtual machine for early microcomputers. The platform established one of the first successful portable gaming ecosystems by providing a hardware abstraction layer that enabled games to run across different systems.
B.2 System Architecture
The CHIP-8 architecture consists of:
  • Memory: 4KB total, with programs loaded at address 0x200
  • Registers: 16 8-bit registers (V0-VF), with VF serving as a flag register
  • Display: 64×32 pixel monochrome screen with XOR-based rendering
  • Input: 16-key hexadecimal keypad (0-9, A-F)
  • Timers: 60Hz delay timer and sound timer
  • Audio: Single-tone buzzer
B.3 Instruction Set Highlights
CHIP-8’s 35-instruction set includes specialized gaming primitives:
  • Sprite Drawing (DXYN): XOR-based rendering enabling collision detection
  • Key Input (EX9E, EXA1): Skip instructions based on key state
  • BCD Conversion (FX33): Convert register values to decimal display
  • Memory Operations: Bulk register loading/storing (FX55, FX65)
The XOR-based sprite system is particularly elegant: drawing the same sprite twice erases it, enabling simple animation and automatic collision detection when pixels turn off.
B.4 Font System
CHIP-8 includes built-in 4×5 pixel font data for hexadecimal digits (0-F), stored at addresses 0x050-0x09F. Games reference these fonts for score and text display by setting the index register to the appropriate font location.
Appendix C Game Environment Implementation Details
C.1 Score Detection Methodology
CHIP-8 games store scoring information in arbitrary memory locations using game-specific formats. Our automated detection operates in two phases:
Static Analysis: We analyze ROM structure for common programming patterns, particularly binary-coded decimal (BCD) operations (FX33 instruction) that suggest numeric display routines.
Dynamic Monitoring: During human gameplay sessions, we monitor memory changes to correlate locations with scoring events. Register trend analysis identifies increasing values (likely scores) versus decreasing values (likely lives/health).
C.2 Game List
C.2.1 Long-horizon Planning & Spatial Reasoning
Requires strategic thinking, spatial awareness, and multi-step planning
  • tetris – Tetris by Fran Dachille (1991): Classic Tetris with piece rotation, movement, and dropping
  • blinky – Blinky by Hans Christian Egeberg (1991): Pac-Man clone with 2 lives, maze with energy pills and 2 ghosts
  • worm – SuperWorm V4 by RB-Revival Studios (2007): Snake-like game with enhanced controls and speed fixes
C.2.2 Timing, Prediction & Reactive Control
Requires precise timing, trajectory prediction, and fast reactive responses
  • brix – Brix by Andreas Gustafsson (1990): Breakout clone with paddle controlling the ball to destroy bricks, 5 lives
  • pong – Pong: Single player pong game with paddle control
  • squash – Squash by David Winter (1997): Bounce ball around squash court with paddle
  • vertical_brix – Vertical Brix by Paul Robson (1996): Breakout variant with vertical brick layout and paddle movement
  • wipe_off – Wipe Off by Joseph Weisbecker: Move paddle to wipe out spots, 1 point per spot, 20 balls
  • filter – Filter: Catch drops from pipe with paddle
C.2.3 Resource Management & Tactical Decisions
Requires managing limited resources and making strategic tactical choices
  • missile – Missile Command by David Winter (1996): Shoot 8 targets using key 8, earn 5 points per target, 12 missiles total
  • rocket – Rocket by Joseph Weisbecker (1978): Launch rockets to hit moving UFO across top of screen, 9 rockets total
  • submarine – Submarine by Carmelo Cortez (1978): Fire depth charges at submarines, 15 points for small, 5 for large subs
  • tank – Tank Battle: Tank with 25 bombs to hit mobile target, lose 5 bombs if tank hits target
  • ufo – UFO by Lutz V (1992): Stationary launcher shoots in 3 directions at flying objects, 15 missiles
C.2.4 Exploration & Continuous Navigation
Requires spatial exploration, obstacle avoidance, and continuous movement control
  • cavern – Cavern by Matthew Mikolay (2014): Navigate cave without hitting walls, modified for leftward exploration
  • flight_runner – Flight Runner by TodPunk (2014): Simple flight navigation game
  • space_flight – Space Flight by Unknown (19xx): Fly through asteroid field using ship navigation controls
  • spacejam – Spacejam! by William Donnelly (2015): Enhanced ship tunnel navigation game
C.2.5 Simple Reaction & Timing
Requires basic reaction time and simple decision making
  • airplane – Airplane: Bombing game where you drop bombs by pressing key 8
  • deep – Deep8 by John Earnest (2014): Move boat left/right, drop and detonate bombs to destroy incoming squid
  • shooting_stars – Shooting Stars by Philip Baltzer (1978): Classic shooting game
Appendix D Hyperparameter Optimization Results
We conducted a comprehensive grid search on the Pong environment to identify optimal PPO hyperparameters before evaluating across the full game suite. The search explored four key dimensions: number of parallel environments, rollout length, minibatch size, and learning rate. All experiments used 4 epochs per update, GAE lambda of 0.95, and gradient clipping at 0.5.
D.1 Search Space
The hyperparameter search explored the following ranges:
  • Environments: $\{128,256,512,1024\}$
  • Rollout steps: $\{32,64,128,512\}$
  • Minibatches: $\{4,8,16,32\}$
  • Learning rate: $\{2.5\times 10^{-4},5\times 10^{-4},1\times 10^{-3}\}$
Each configuration was trained for 1M timesteps with evaluation every 65,536 steps. Final evaluation scores represent the last recorded performance, where less negative values indicate better performance.
D.2 Results Summary
Table 2 presents the key configurations and their final evaluation scores. Higher scores indicate better performance (scores are negative, with values closer to zero being better).
Table 2: Hyperparameter search results on Pong environment. Configurations sorted by final evaluation score.
Envs Steps Minibatches LR Score
512 32 32 0.0005 -2.34
512 32 16 0.001 -2.48
512 32 32 0.001 -2.69
128 128 16 0.00025 -2.95
128 64 8 0.00025 -3.19
512 32 16 0.00025 -3.20
256 64 16 0.00025 -3.38
128 32 4 0.00025 -3.44
128 64 16 0.00025 -3.53
512 32 16 0.0005 -3.73
256 32 4 0.00025 -3.78
128 128 8 0.00025 -3.91
256 128 32 0.00025 -4.03
512 64 16 0.00025 -4.17
1024 32 32 0.00025 -4.34
1024 32 16 0.00025 -4.44
256 128 16 0.00025 -4.66
1024 64 32 0.00025 -4.96
D.3 Analysis and Key Findings
Learning rate impact. Higher learning rates significantly improved performance, with $5\times 10^{-4}$ and $1\times 10^{-3}$ substantially outperforming $2.5\times 10^{-4}$. The top three configurations all used learning rates above the commonly used $2.5\times 10^{-4}$.
Environment scaling. 512 parallel environments provided the optimal balance between computational efficiency and sample diversity. Configurations with 1024 environments showed diminishing returns, possibly due to computational overhead or reduced gradient update frequency.
Rollout length. Shorter rollouts (32 steps) consistently outperformed longer ones, indicating more frequent policy updates may be beneficial for this environment.
Minibatch size. Larger minibatch sizes (16-32) generally improved performance by providing more stable gradient estimates, though the effect was less pronounced than learning rate changes.
D.4 Final Configuration
Based on these results, we selected the following hyperparameters for all subsequent experiments:
  • Parallel environments: 512
  • Rollout steps: 32
  • Training epochs: 4
  • Minibatches: 32
  • Learning rate: $5\times 10^{-4}$
  • GAE lambda: 0.95
  • Clip epsilon: 0.2
  • Value function coefficient: 0.5
  • Entropy coefficient: 0.01
Appendix E LLM-Assisted Environment Generation
This appendix details the automated environment generation pipeline using large language models (LLMs) to create novel CHIP-8 games for reinforcement learning research. We demonstrate the complete process from prompt engineering to code generation across three difficulty levels of a Target Shooter game.
E.1 Prompt Engineering
Our LLM generation pipeline relies on carefully crafted prompts that provide comprehensive CHIP-8 programming context and specific game requirements. The core prompt structure includes CHIP-8 architectural constraints, Octo assembly language syntax, and reinforcement learning compatibility requirements.
Listing 1: LLM prompt template for CHIP-8 game generation
You are a **professional CHIP-8 (classic version) game developer**.
Your task is to **design and implement new CHIP-8 games in Octo assembly language**. I will provide you with tutorials and references for Octo assembly. You must be rigorous and ensure that your code is **syntactically correct, runnable, and follows CHIP-8 conventions**.
<documentation></documentation>
<tutorial1></tutorial1>
<tutorial2></tutorial2>
<example></example>
The goal is to create a **game suitable for reinforcement learning (RL) research**, which means:
* The **score** must be stored in a clear and consistent register or memory location.
* The **termination condition** (game over) must also be easily extractable (e.g., through a specific flag or register value).
* The game should have **deterministic rules** and be lightweight enough for training agents.
Here is the description of the game you must implement:
<description>{{description}}</description>
The prompt incorporates several key components:
  • Role specification: Establishes the LLM as a professional CHIP-8 developer
  • Technical constraints: Emphasizes syntactic correctness and CHIP-8 compliance
  • RL compatibility: Specifies requirements for score tracking and termination detection
  • Reference material: Includes comprehensive CHIP-8 documentation and examples
  • Game description: Placeholder for specific game mechanics and objectives
The prompt template includes placeholder tags that are populated with comprehensive CHIP-8 resources: <documentation> contains the official Octo Manual (https://johnearnest.github.io/Octo/docs/Manual.html), <tutorial1> includes the Beginner’s Guide (https://johnearnest.github.io/Octo/docs/BeginnersGuide.html), <tutorial2> incorporates the Intermediate Guide (https://johnearnest.github.io/Octo/docs/IntermediateGuide.html), and <example> provides a complete game implementation (https://github.com/JohnEarnest/chip8Archive/blob/master/src/outlaw/outlaw.8o) to demonstrate best practices and coding patterns.
E.2 Generated Target Shooter Implementation
Using the prompt template, we generated three progressive difficulty levels of a Target Shooter game. Each level maintains consistent register mappings for score and termination while introducing increasing complexity in target behavior and timing constraints.
E.2.1 Level 1: Static Targets
The first difficulty level features stationary targets that appear at random locations, focusing on basic aiming and shooting mechanics.
Listing 2: Level 1 Target Shooter - Static targets
# Target Shooter - RL Training Game
# A deterministic shooting game designed for
# reinforcement learning research.
# Controls:
# - WASD to move crosshair
# - E to shoot
# Score is stored in register v2 (score_reg)
# Game over flag in register v3 (gameover_reg)
# Game ends after hitting 10 targets.
# Sprite data
: crosshair
0b10000001
0b01011010
0b00100100
0b01011010
0b01011010
0b00100100
0b01011010
0b10000001
: target
0b00111100
0b01000010
0b10011001
0b10100101
0b10100101
0b10011001
0b01000010
0b00111100
# Register Map - Critical for RL extraction
:alias crosshair_x v0 # Crosshair X position
:alias crosshair_y v1 # Crosshair Y position
:alias score_reg v2 # SCORE - RL agents read this!
:alias gameover_reg v3 # GAME OVER FLAG (0=playing, 1=over)
:alias target_x v4 # Target X position
:alias target_y v5 # Target Y position
:alias target_active v6 # Target active flag
:alias temp1 v7 # Temporary register
:alias temp2 v8 # Temporary register
:alias shot_active v9 # Shot in progress flag
:alias targets_hit va # Count of targets hit (max 10)
:alias key_reg vb # Key input register
:const MAX_TARGETS 10 # Game ends after 10 targets
:const TARGET_SIZE 8 # Target sprite size
:const CROSSHAIR_SIZE 8 # Crosshair sprite size
:const POINTS_PER_HIT 1 # Points awarded per target hit
# Main Game Entry Point
: main
# Initialize game state
score_reg := 0 # Score starts at 0
gameover_reg := 0 # Game is not over
targets_hit := 0 # No targets hit yet
target_active := 0 # No target active initially
shot_active := 0 # No shot in progress
# Initial crosshair position (center)
crosshair_x := 28
crosshair_y := 12
clear
# Draw initial UI
draw-crosshair
# Main game loop
loop
# Check if game should end
if targets_hit == MAX_TARGETS then jump game-over
# Spawn new target if none active
if target_active == 0 then spawn-target
# Handle player input
handle-input
# Check for hit if shot was fired
if shot_active == 1 then check-hit
# Small delay for playability
temp1 := 1
delay := temp1
wait-delay
again
# Game Over Handler
: game-over
gameover_reg := 1 # Set game over flag for RL agent
# Flash screen to indicate game over
temp1 := 0
loop
clear
temp2 := 5
delay := temp2
wait-delay
draw-crosshair
if target_active == 1 then draw-target
temp2 := 5
delay := temp2
wait-delay
temp1 += 1
if temp1 != 3 then
again
# Infinite loop - game is over
loop
# RL agent should detect gameover_reg == 1
again
# Input Handling
: handle-input
# Save current position
temp1 := crosshair_x
temp2 := crosshair_y
# Movement controls (WASD) - use consistent key codes
key_reg := 7 # A key (left)
if key_reg key then temp1 += -2
key_reg := 9 # D key (right)
if key_reg key then temp1 += 2
key_reg := 5 # W key (up)
if key_reg key then temp2 += -2
key_reg := 8 # S key (down)
if key_reg key then temp2 += 2
# Boundary checking
if temp1 >= 254 then temp1 := 0 # Left boundary (wrapping check)
if temp1 >= 56 then temp1 := 56 # Right boundary
if temp2 >= 254 then temp2 := 0 # Top boundary (wrapping check)
if temp2 >= 24 then temp2 := 24 # Bottom boundary
# Check if position changed
if temp1 != crosshair_x then jump update-crosshair
if temp2 != crosshair_y then jump update-crosshair
# Check for shoot (E key)
key_reg := 6
if key_reg key then shot_active := 1
return
: update-crosshair
# Erase old crosshair
i := crosshair
sprite crosshair_x crosshair_y CROSSHAIR_SIZE
# Update position
crosshair_x := temp1
crosshair_y := temp2
# Draw new crosshair
sprite crosshair_x crosshair_y CROSSHAIR_SIZE
# Check shoot after movement
key_reg := 6
if key_reg key then shot_active := 1
# Target Management
: spawn-target
# Generate random position for target
target_x := random 0x37 # 0-55 range
target_y := random 0x17 # 0-23 range
# Ensure minimum distance from edges
if target_x <= 2 then target_x := 3
if target_y <= 2 then target_y := 3
target_active := 1
draw-target
: draw-target
i := target
sprite target_x target_y TARGET_SIZE
: draw-crosshair
i := crosshair
sprite crosshair_x crosshair_y CROSSHAIR_SIZE
# Hit Detection
: check-hit
shot_active := 0 # Reset shot flag
# Check if target is active
if target_active == 0 then return
# Simple hit detection - check if crosshair center is near target center
# Calculate X distance
temp1 := crosshair_x
temp1 += 4 # Crosshair center
temp2 := target_x
temp2 += 4 # Target center
# Check X proximity
if temp1 > temp2 then jump check-x-greater
# crosshair is left of or at target
temp2 -= temp1
if temp2 > 6 then return # Too far
jump check-y-axis
: check-x-greater
# crosshair is right of target
temp1 -= temp2
if temp1 > 6 then return # Too far
: check-y-axis
# Calculate Y distance
temp1 := crosshair_y
temp1 += 4 # Crosshair center
temp2 := target_y
temp2 += 4 # Target center
# Check Y proximity
if temp1 > temp2 then jump check-y-greater
# crosshair is above or at target
temp2 -= temp1
if temp2 > 6 then return # Too far
jump register-hit
: check-y-greater
# crosshair is below target
temp1 -= temp2
if temp1 > 6 then return # Too far
: register-hit
# Hit confirmed!
# Erase target
draw-target
target_active := 0
# Update score (for RL agent)
score_reg += POINTS_PER_HIT
targets_hit += 1
# Sound feedback
temp1 := 3
buzzer := temp1
# Utility Functions
: wait-delay
loop
temp1 := delay
if temp1 != 0 then
again
E.2.2 Level 2: Time-Limited Targets
The second level introduces time pressure by making targets disappear after a fixed duration, requiring faster decision-making from RL agents.
Listing 3: Level 2 Target Shooter - Time-limited targets
# Target Shooter - RL Training Game
# A deterministic shooting game designed for
# reinforcement learning research.
# Controls:
# - WASD to move crosshair
# - E to shoot
# Score is stored in register v2 (score_reg)
# Game over flag in register v3 (gameover_reg)
# Game ends after 10 targets (hit or missed).
# Targets disappear after ~3 seconds if not hit.
# Sprite data
: crosshair
0b10000001
0b01011010
0b00100100
0b01011010
0b01011010
0b00100100
0b01011010
0b10000001
: target
0b00111100
0b01000010
0b10011001
0b10100101
0b10100101
0b10011001
0b01000010
0b00111100
# Register Map - Critical for RL extraction
:alias crosshair_x v0 # Crosshair X position
:alias crosshair_y v1 # Crosshair Y position
:alias score_reg v2 # SCORE - RL agents read this!
:alias gameover_reg v3 # GAME OVER FLAG (0=playing, 1=over)
:alias target_x v4 # Target X position
:alias target_y v5 # Target Y position
:alias target_active v6 # Target active flag
:alias temp1 v7 # Temporary register
:alias temp2 v8 # Temporary register
:alias shot_active v9 # Shot in progress flag
:alias targets_total va # Total targets appeared (max 10)
:alias key_reg vb # Key input register
:alias target_timer vc # Timer for current target
:alias missed_targets vd # Count of missed targets
:const MAX_TARGETS 10 # Game ends after 10 targets total
:const TARGET_SIZE 8 # Target sprite size
:const CROSSHAIR_SIZE 8 # Crosshair sprite size
:const POINTS_PER_HIT 1 # Points awarded per target hit
:const TARGET_TIMEOUT 60 # Frames before target disappears (~3 sec at 20fps)
# Main Game Entry Point
: main
# Initialize game state
score_reg := 0 # Score starts at 0
gameover_reg := 0 # Game is not over
targets_total := 0 # No targets appeared yet
missed_targets := 0 # No missed targets yet
target_active := 0 # No target active initially
shot_active := 0 # No shot in progress
target_timer := 0 # Timer at 0
# Initial crosshair position (center)
crosshair_x := 28
crosshair_y := 12
clear
# Draw initial UI
draw-crosshair
# Main game loop
loop
# Check if game should end (10 total targets)
if targets_total == MAX_TARGETS then jump game-over
# Spawn new target if none active
if target_active == 0 then spawn-target
# Check target timeout
if target_active == 1 then check-target-timeout
# Handle player input
handle-input
# Check for hit if shot was fired
if shot_active == 1 then check-hit
# Small delay for playability
temp1 := 1
delay := temp1
wait-delay
again
# Target Timeout Check
: check-target-timeout
# Decrement timer
target_timer += -1
# Check if timer expired
if target_timer != 0 then return
# Target timed out - count as miss
draw-target # Erase target
target_active := 0
missed_targets += 1
# Brief sound to indicate miss
temp1 := 1
buzzer := temp1
# Game Over Handler
: game-over
gameover_reg := 1 # Set game over flag for RL agent
# Flash screen to indicate game over
temp1 := 0
loop
clear
temp2 := 5
delay := temp2
wait-delay
draw-crosshair
if target_active == 1 then draw-target
temp2 := 5
delay := temp2
wait-delay
temp1 += 1
if temp1 != 3 then
again
# Infinite loop - game is over
loop
# RL agent should detect gameover_reg == 1
again
# Input Handling
: handle-input
# Save current position
temp1 := crosshair_x
temp2 := crosshair_y
# Movement controls (WASD) - use consistent key codes
key_reg := 7 # A key (left)
if key_reg key then temp1 += -2
key_reg := 9 # D key (right)
if key_reg key then temp1 += 2
key_reg := 5 # W key (up)
if key_reg key then temp2 += -2
key_reg := 8 # S key (down)
if key_reg key then temp2 += 2
# Boundary checking
if temp1 >= 254 then temp1 := 0 # Left boundary (wrapping check)
if temp1 >= 56 then temp1 := 56 # Right boundary
if temp2 >= 254 then temp2 := 0 # Top boundary (wrapping check)
if temp2 >= 24 then temp2 := 24 # Bottom boundary
# Check if position changed
if temp1 != crosshair_x then jump update-crosshair
if temp2 != crosshair_y then jump update-crosshair
# Check for shoot (E key)
key_reg := 6
if key_reg key then shot_active := 1
return
: update-crosshair
# Erase old crosshair
i := crosshair
sprite crosshair_x crosshair_y CROSSHAIR_SIZE
# Update position
crosshair_x := temp1
crosshair_y := temp2
# Draw new crosshair
sprite crosshair_x crosshair_y CROSSHAIR_SIZE
# Check shoot after movement
key_reg := 6
if key_reg key then shot_active := 1
# Target Management
: spawn-target
# Generate random position for target
target_x := random 0x37 # 0-55 range
target_y := random 0x17 # 0-23 range
# Ensure minimum distance from edges
if target_x <= 2 then target_x := 3
if target_y <= 2 then target_y := 3
target_active := 1
target_timer := TARGET_TIMEOUT # Set timeout timer
targets_total += 1 # Increment total targets count
draw-target
: draw-target
i := target
sprite target_x target_y TARGET_SIZE
: draw-crosshair
i := crosshair
sprite crosshair_x crosshair_y CROSSHAIR_SIZE
# Hit Detection
: check-hit
shot_active := 0 # Reset shot flag
# Check if target is active
if target_active == 0 then return
# Simple hit detection - check if crosshair center is near target center
# Calculate X distance
temp1 := crosshair_x
temp1 += 4 # Crosshair center
temp2 := target_x
temp2 += 4 # Target center
# Check X proximity
if temp1 > temp2 then jump check-x-greater
# crosshair is left of or at target
temp2 -= temp1
if temp2 > 6 then return # Too far
jump check-y-axis
: check-x-greater
# crosshair is right of target
temp1 -= temp2
if temp1 > 6 then return # Too far
: check-y-axis
# Calculate Y distance
temp1 := crosshair_y
temp1 += 4 # Crosshair center
temp2 := target_y
temp2 += 4 # Target center
# Check Y proximity
if temp1 > temp2 then jump check-y-greater
# crosshair is above or at target
temp2 -= temp1
if temp2 > 6 then return # Too far
jump register-hit
: check-y-greater
# crosshair is below target
temp1 -= temp2
if temp1 > 6 then return # Too far
: register-hit
# Hit confirmed!
# Erase target
draw-target
target_active := 0
target_timer := 0 # Clear timer
# Update score (for RL agent)
score_reg += POINTS_PER_HIT
# Sound feedback
temp1 := 3
buzzer := temp1
# Utility Functions
: wait-delay
loop
temp1 := delay
if temp1 != 0 then
again
E.2.3 Level 3: Moving Targets with Time Constraints
The most challenging level combines target movement with time limits, requiring predictive aiming and rapid response times.
Listing 4: Level 3 Target Shooter - Moving targets with time constraints
# Target Shooter Level 3 - RL Training Game
# A deterministic shooting game designed for
# reinforcement learning research.
# Controls:
# - WASD to move crosshair
# - E to shoot
# Score is stored in register v2 (score_reg)
# Game over flag in register v3 (gameover_reg)
# Features:
# - Moving targets that bounce off walls
# - Targets disappear after ~3 seconds if not hit
# - Game ends after 10 targets (hit or missed)
# Sprite data
: crosshair
0b10000001
0b01011010
0b00100100
0b01011010
0b01011010
0b00100100
0b01011010
0b10000001
: target
0b00111100
0b01000010
0b10011001
0b10100101
0b10100101
0b10011001
0b01000010
0b00111100
# Register Map - Critical for RL extraction
:alias crosshair_x v0 # Crosshair X position
:alias crosshair_y v1 # Crosshair Y position
:alias score_reg v2 # SCORE - RL agents read this!
:alias gameover_reg v3 # GAME OVER FLAG (0=playing, 1=over)
:alias target_x v4 # Target X position
:alias target_y v5 # Target Y position
:alias target_active v6 # Target active flag
:alias temp1 v7 # Temporary register
:alias temp2 v8 # Temporary register
:alias shot_active v9 # Shot in progress flag
:alias targets_total va # Total targets appeared (max 10)
:alias key_reg vb # Key input register
:alias target_timer vc # Timer for current target
:alias target_vx vd # Target X velocity
:alias target_vy ve # Target Y velocity
:const MAX_TARGETS 10 # Game ends after 10 targets total
:const TARGET_SIZE 8 # Target sprite size
:const CROSSHAIR_SIZE 8 # Crosshair sprite size
:const POINTS_PER_HIT 1 # Points awarded per target hit
:const TARGET_TIMEOUT 80 # Frames before target disappears (~4 sec with movement)
# Main Game Entry Point
: main
# Initialize game state
score_reg := 0 # Score starts at 0
gameover_reg := 0 # Game is not over
targets_total := 0 # No targets appeared yet
target_active := 0 # No target active initially
shot_active := 0 # No shot in progress
target_timer := 0 # Timer at 0
target_vx := 0 # No initial velocity
target_vy := 0 # No initial velocity
# Initial crosshair position (center)
crosshair_x := 28
crosshair_y := 12
clear
# Draw initial UI
draw-crosshair
# Main game loop
loop
# Check if game should end (10 total targets)
if targets_total == MAX_TARGETS then jump game-over
# Spawn new target if none active
if target_active == 0 then spawn-target
# Update target position if active
if target_active == 1 then move-target
# Check target timeout
if target_active == 1 then check-target-timeout
# Handle player input
handle-input
# Check for hit if shot was fired
if shot_active == 1 then check-hit
# Small delay for playability
temp1 := 1
delay := temp1
wait-delay
again
# Target Movement
: move-target
# Erase target at current position
draw-target
# Update X position
target_x += target_vx
# Check X boundaries and bounce
if target_x >= 250 then jump bounce-left # Hit left edge
if target_x >= 56 then jump bounce-right # Hit right edge
: check-y-movement
# Update Y position
target_y += target_vy
# Check Y boundaries and bounce
if target_y >= 250 then jump bounce-top # Hit top edge
if target_y >= 24 then jump bounce-bottom # Hit bottom edge
: finish-move
# Draw target at new position
draw-target
return
: bounce-left
target_x := 1
target_vx := 1 # Reverse to move right
jump check-y-movement
: bounce-right
target_x := 55
target_vx := 255 # -1 to move left
jump check-y-movement
: bounce-top
target_y := 1
target_vy := 1 # Reverse to move down
jump finish-move
: bounce-bottom
target_y := 23
target_vy := 255 # -1 to move up
jump finish-move
# Target Timeout Check
: check-target-timeout
# Decrement timer
target_timer += -1
# Check if timer expired
if target_timer != 0 then return
# Target timed out - count as miss
draw-target # Erase target
target_active := 0
# Brief sound to indicate miss
temp1 := 1
buzzer := temp1
# Game Over Handler
: game-over
gameover_reg := 1 # Set game over flag for RL agent
# Flash screen to indicate game over
temp1 := 0
loop
clear
temp2 := 5
delay := temp2
wait-delay
draw-crosshair
if target_active == 1 then draw-target
temp2 := 5
delay := temp2
wait-delay
temp1 += 1
if temp1 != 3 then
again
# Infinite loop - game is over
loop
# RL agent should detect gameover_reg == 1
again
# Input Handling
: handle-input
# Save current position
temp1 := crosshair_x
temp2 := crosshair_y
# Movement controls (WASD) - use consistent key codes
key_reg := 7 # A key (left)
if key_reg key then temp1 += -2
key_reg := 9 # D key (right)
if key_reg key then temp1 += 2
key_reg := 5 # W key (up)
if key_reg key then temp2 += -2
key_reg := 8 # S key (down)
if key_reg key then temp2 += 2
# Boundary checking
if temp1 >= 254 then temp1 := 0 # Left boundary (wrapping check)
if temp1 >= 56 then temp1 := 56 # Right boundary
if temp2 >= 254 then temp2 := 0 # Top boundary (wrapping check)
if temp2 >= 24 then temp2 := 24 # Bottom boundary
# Check if position changed
if temp1 != crosshair_x then jump update-crosshair
if temp2 != crosshair_y then jump update-crosshair
# Check for shoot (E key)
key_reg := 6
if key_reg key then shot_active := 1
return
: update-crosshair
# Erase old crosshair
i := crosshair
sprite crosshair_x crosshair_y CROSSHAIR_SIZE
# Update position
crosshair_x := temp1
crosshair_y := temp2
# Draw new crosshair
sprite crosshair_x crosshair_y CROSSHAIR_SIZE
# Check shoot after movement
key_reg := 6
if key_reg key then shot_active := 1
# Target Management
: spawn-target
# Generate random position for target
target_x := random 0x37 # 0-55 range
target_y := random 0x17 # 0-23 range
# Ensure minimum distance from edges
if target_x <= 2 then target_x := 3
if target_y <= 2 then target_y := 3
# Generate random velocity (-1, 0, or 1 for each axis)
target_vx := random 0x03
if target_vx == 2 then target_vx := 255 # Convert 2 to -1
target_vy := random 0x03
if target_vy == 2 then target_vy := 255 # Convert 2 to -1
# Ensure target is moving (not both velocities zero)
if target_vx == 0 then jump ensure-movement
jump finish-spawn
: ensure-movement
if target_vy == 0 then target_vy := 1
: finish-spawn
target_active := 1
target_timer := TARGET_TIMEOUT # Set timeout timer
targets_total += 1 # Increment total targets count
draw-target
: draw-target
i := target
sprite target_x target_y TARGET_SIZE
: draw-crosshair
i := crosshair
sprite crosshair_x crosshair_y CROSSHAIR_SIZE
# Hit Detection
: check-hit
shot_active := 0 # Reset shot flag
# Check if target is active
if target_active == 0 then return
# Simple hit detection - check if crosshair center is near target center
# Calculate X distance
temp1 := crosshair_x
temp1 += 4 # Crosshair center
temp2 := target_x
temp2 += 4 # Target center
# Check X proximity
if temp1 > temp2 then jump check-x-greater
# crosshair is left of or at target
temp2 -= temp1
if temp2 > 6 then return # Too far
jump check-y-axis
: check-x-greater
# crosshair is right of target
temp1 -= temp2
if temp1 > 6 then return # Too far
: check-y-axis
# Calculate Y distance
temp1 := crosshair_y
temp1 += 4 # Crosshair center
temp2 := target_y
temp2 += 4 # Target center
# Check Y proximity
if temp1 > temp2 then jump check-y-greater
# crosshair is above or at target
temp2 -= temp1
if temp2 > 6 then return # Too far
jump register-hit
: check-y-greater
# crosshair is below target
temp1 -= temp2
if temp1 > 6 then return # Too far
: register-hit
# Hit confirmed!
# Erase target
draw-target
target_active := 0
target_timer := 0 # Clear timer
# Update score (for RL agent)
score_reg += POINTS_PER_HIT
# Sound feedback
temp1 := 3
buzzer := temp1
# Utility Functions
: wait-delay
loop
temp1 := delay
if temp1 != 0 then
again
E.2.4 Environment Integration and Wrapper Implementation
Once the LLM generates the CHIP-8 assembly code for each difficulty level, the games require integration with Octax’s reinforcement learning interface. The environment wrapper extracts reward signals and termination conditions from the consistent register mapping established during code generation.
The Target Shooter implementation demonstrates the integration between LLM-generated content and the Octax framework. Each level maintains identical register assignments to ensure compatibility across the difficulty progression, enabling curriculum learning experiments without code modifications.
Listing 5: Target Shooter environment wrapper implementation
from octax import EmulatorState
def score_fn(state: EmulatorState) -> float:
␣␣␣␣ExtractscorefromregisterV[2]
␣␣␣␣Scoreincrementsby1foreachsuccessfulhit
␣␣␣␣Range:0-10points
return state.V[2]
def terminated_fn(state: EmulatorState) -> bool:
␣␣␣␣CheckgameterminationflaginregisterV[3]
␣␣␣␣Gameendsafter10totaltargets(hitormissedinlevels2-3)
return state.V[3] == 1
# CHIP-8 key mapping for controls
# W=5 (up), A=7 (left), S=8 (down), D=9 (right), E=6 (shoot)
action_set = [5, 7, 8, 9, 6]
metadata = {
"title": "TargetShooter-LLM-GeneratedRLEnvironment",
"authors": ["FullyLLM-GeneratedEnvironment"],
"description": "AI-generatedprogressivedifficultyenvironment",
"roms": {
"target_shooter_level1": {
"file": "target_shooter_level1.ch8",
"description": "Statictargets-Basicaimingskills"
"target_shooter_level2": {
"file": "target_shooter_level2.ch8",
"description": "Time-limitedstatictargets"
"target_shooter_level3": {
"file": "target_shooter_level3.ch8",
"description": "Movingtime-limitedtargets"
The consistent register mapping across all three levels enables direct comparison of agent performance and facilitates automated curriculum progression. Register V[2] consistently stores the score for reward calculation, while V[3] serves as the binary termination flag. The five-action control scheme (WASD movement plus shoot) provides sufficient complexity for interesting policies while remaining tractable for systematic analysis.